OpenStudy (lukecrayonz):

e^(2x)-15e^x+56=0 e^(2x)-15e^x+56=0 @Mathematics

OpenStudy (anonymous):

solve \[x^2-15x+56=0\] first and then replace x by e^x

OpenStudy (anonymous):

or just factor as \[(e^x-7)(e^x-8)=0\] \[e^x=7, e^x=8\] \[x=\ln(7), x=\ln(8)\]